OpenStudy (anonymous):

HELP WITH CIRCLES! Find equations for the circle and tangent line: Center of the circle (0,0). Line tangent to circle at (-5,12).

OpenStudy (anonymous):

x^2+y^2=13 :) tangent is xx1+yy1=0

OpenStudy (anonymous):

how do you find this? I don't need just the answer, i need to know the steps to solve please

OpenStudy (noelgreco):

Draw a circle centered at the origin. Plot a point at (-5,12). Draw a radius from the origin to the point. That gives you a right triangle, which turns out to be a Pythagorean triple giving you the radius. Then:\[x ^{2}+y ^{2}=r ^{2}\]

OpenStudy (anonymous):

 and (x-x1)^2+(y-y1)^2=r^2:)

OpenStudy (anonymous):

so the equation of the circle is x^2+y^2=169?